OpenStudy (gabbyalicorn):

What is the approximate volume of a sphere with a radius of 9 in.? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (gabbyalicorn):

@kohai

OpenStudy (triciaal):

use the same formula I just gave you. this time substitute for pi and you already have r

OpenStudy (gabbyalicorn):

4/3 x 3.14 x 81

OpenStudy (triciaal):



OpenStudy (anonymous):

\[\frac{4 \pi 9^3}{3}=972 \pi =3053.63 \text{ cubic inches}\]

OpenStudy (gabbyalicorn):

so 4 x 3.14 x ÷ 3

OpenStudy (anonymous):

( 4*3.14*9*9*9 )/3 = (9156.24)/3

OpenStudy (gabbyalicorn):

i got 3052.08 is this correct? @robtobey

OpenStudy (gabbyalicorn):

yep it is. :) ty

OpenStudy (gabbyalicorn):

also if i was doing this on another problem could i use 4*3.14*9*9*9

OpenStudy (gabbyalicorn):

but how'd you get to the third power. i read the problem and I didn't see it say it has a 3.